A report says that the average amount of time a 10-year-old American child spends playing outdoors per day is between 20.32 and 24.56 minutes. What is the margin of error in this report?
Answer #1

\bar{x}=\frac{20.32+24.56}{2}=22.44

E=24.56-22.44=2.12

Margin of error =2.12